OpenStudy (anonymous):

What is the radical expression of a^5/7?

OpenStudy (anonymous):

example: \[\sqrt[2]{a ^{1}} = \sqrt{a}\]

OpenStudy (anonymous):

\[\sqrt{a}=a ^{1/2}\]

OpenStudy (anonymous):

So, \[\sqrt[7]{a ^{5}}\]

OpenStudy (anonymous):

\[\sqrt{\frac{a^{5/2}}{7}}=\frac{a^{5/2}\sqrt{7}}{7} \]